Presented below is information related to Splish Inc SPLISH INC BALANCE SHEET DECEMBER 31, 2017 Cash $44,600 Notes payable (short-term) Accounts payable $50,100 32,200 4,900 258,400 139,700 Receivables $109,700 15,000 94,700 Accrued liabilities 169,200 Common stock (par $5) Less: Allowance Inventory Prepaid insurance Land Equipment (net) 8,000 Retained earnings 20,000 148,800 $485,300 $485,300 SPLISH INC INCOME STATEMENT FOR THE YEAR ENDED DECEMBER 31, 2017 Sales revenue $1,395,000 Cost of goods sold Inventory, Jan. 1, 2017 Purchases Cost of goods available for sale Inventory, Dec. 31, 2017 Cost of goods sold $198,600 792,200 990,800 (169,200) Gross profit on sales Operating expenses Net income 821,600 573,400 171,500 $401,900 Compute the following ratios or relationships of Splish Inc. Assume that the ending account balances are representative unless the information provided indicates differently. (Round answers to 2 decimal places, e.g. 1.59 or 45.87%.) (1) Current ratio (2) Inventory turnover (3) Accounts receivable turnover (4) Earnings per share (5) Profit margin on sales (6) Return on assets on December 31, 2017 times times times
Answer #1
1)Current ratio
Current assets=44600+94700+169200+8000= 316500
Current liabilities=50100+32200+4900= 87200
Current ratio = 316500/87200=           3.63
2)Inventory turnover
Cost of goods sold 821600
Average inventory (198600+169200)/2 183900
Inventory turnover =COGS/avg inventory           4.47
3)Receivables turnover
Sales/Accounts receivable=1395000/94700         14.73
4)Earnings per share
Net income 401900
No. of shares(258400/$5) 51680
EPS =NET INCOME/no. of shares
(401900/51680) $       7.78
5)Profit margin on sales
Net income/sales 28.81%
(401900/1395000)
6)Return on assets
Net income/total assets 82.81%
(401900/485300)